OpenStudy (anonymous):

To buy both a new car and a new house, Tina sought two loans totalling $318,338. The simple interest rate on the first loan was 2.4%, while the simple interest rate on the second loan was 4.7%. At the end of the first year, Tina paid a combined interest payment of $14,761.81. What were the amounts of the two loans?

OpenStudy (anonymous):

0.024x+0.047(318338-x)=14761 0.024x+14961.89-0.047x=14761 -0.023x=-200 x=$8,695.65 at 2.4% 318338-x=$309,642.35 at 4.7%

OpenStudy (anonymous):

is this right

OpenStudy (lalaly):

yeh thats right

OpenStudy (anonymous):

ty